# Functional Protein Design with Local Domain Alignment

### v2

@@ 3.4 Annotation-guided Protein Design
Definition 3.1 (Annotation-guided Protein Design)
@@
$\vspace{-1ex}\max\sum_{T\in\mathcal{A}}\mathrm{M}_{T}(S),\text{ s.t. }S=F(\mathcal{A}).$ (10)
where $\mathrm{M}_{T}()$ is corresponding metric function for the annotation $T$.
In this paper, we mainly focus on two type of measures, functional-domain measure and global-property measure.
  • Functional-domain metric: for the annotation describing a certain functional domain, $\mathrm{M}_{T}()$ will invoke a profile hidden Markov models from Pfam Mistry et al. (2021) to search for the optimal match within protein $S$ regarding to the domain described by annotation $T$ and assign an e-value $s$ to this match. Given an e-value threshold $e$, we have:
    $\mathrm{M}_{T,e}(S)=\mathbbm{1}_{S}(s<e),$ (11)
@@ 4 Experiment
4.1 Construction of ProtAnnotation Dataset
To enable multi-level alignment, we build the ProtAnnotation dataset with annotation-sequence pair set for protein design task. Specifically, we select all proteins with “Domain” entry in UniProtKB Consortium (2022) to build ProtAnnotation dataset, resulting $129,727$ proteins. The domain annotations are extracted by these “Domain” entries including the domain description and start & end index of this domain. Additionally, we select four properties as the property annotations, that is, “protein_name”, “organism_name”, “length” and “SIMLARITY”. The textual description is assembled by the template function $G$, which can accept any subset of annotation as input. More details, including the data examples are deferred in Appendix A.2.
4.2 Quality of Aligned Representation
We first conduct multiple experiments on predictive tasks. to evaluate the quality of protein representation produced by PAAG.
Settings To make a fair comparison with ProtST, we use the same proteins in dataset ProtDescribe Xu et al. (2023) to first pretrain PAAG, followed by full-model fine-tuning on various downstream tasks. For the full-model fine-tuning, we add a task head for each task and fine-tune the model for 100 epochs. We use the validation set to select the model and report the results on random seed 0, adhering to the same settings as in ProtST Xu et al. (2023). More details are deferred in Appendix A.6.
@@ 4.2 Quality of Aligned Representation
  • Protein Localization Prediction aims to forecast the subcellular locations of proteins. Derived from DeepLoc Almagro Armenteros et al. (2017), we focus on two similar tasks. The subcellular localization prediction (Abbr. as, Sub) encompasses 10 location categories and binary localization prediction (Abbr. as, Bin) that includes 2 location categories, soluble and membrane-bound. The splits of data follow the original split in DeepLoc.
  • Fitness Landscape Prediction is primarily focused on the prediction of the effects of residue mutations on the fitness of proteins. We evaluate our models on $\beta$-lactamase (Abbr., $\beta$-lac) landscape from PEER Xu et al. (2022), the AAV and Thermostability (Abbr., Thermo) landscapes from FLIP Dallago et al. (2021), and the Fluorescence (Abbr., Flu) and Stability (Abbr., Sta) landscapes from TAPE Rao et al. (2019). The splits of data follow the splitting setting of ProtST Xu et al. (2023)
Baselines: We adopt two types of baseline. The first type is the models train from scratch, including CNN Shanehsazzadeh et al. (2020), ResNet Rao et al. (2019), LSTM Rao et al. (2019), Transformer Rao et al. (2019). The second type is the pretrained models with full model tuning, including OntoProtein Zhang et al. (2022), ProtBert Elnaggar et al. (2021) and ESM2 Lin et al. (2023). For the ProtST and PAAG, we train two variants with different initialization weights from ProtBert and ESM2 to verify the enhancement of text knowledge on protein representations from different protein encoders. We utilize distinct subscripts to denote the initialized parameters, such as $\text{PAAG}_{\text{ProtBert}}$ and $\text{PAAG}_{\text{ESM2}}$.
Table 1: Results on protein localization prediction and protein landscape prediction benchmarks. Bold in green and underlined numbers indicate the best and the second best result, respectively.
@@
$\text{ProtST}_{\text{ProtBert}}$ 91.78 78.71 0.863 0.804 0.673 0.679 0.745
$\text{ProtST}_{\text{ESM2}}$ 92.52 80.22 0.879 0.825 0.682 0.682 0.738
$\text{PAAG}_{\text{ProtBert}}$ 92.63 78.96 0.820 0.825 0.668 0.680 0.788
$\text{PAAG}_{\text{ESM2}}$ 92.46 81.30 0.888 0.839 0.684 0.682 0.737
Figure 3: Visualization of jointly generation with domain and property annotations. PAAG is capable of integrating domain and property annotations.
Results: Table 1 reports the results of all models on seven baselines. As illustrated in Table 1, PAAG achieves superior performance in comparison to the baselines on all 7 tasks. These results highlight the robust generalization capabilities of PAAG for downstream tasks. Specifically, PAAG outperforms the vanilla pretrained models, ProtBert and ESM2, in all cases, indicating that PAAG can further enhance the quality of protein representation by incorporating the knowledge from textual annotations. Furthermore, to assess the impact of multimodal training on enhancing protein representations, we visualize the relative improvements of ProtST and PAAG in comparison to the vanilla pre-trained models in Figure 3. As observed in Figure 3, multi-modal training has a positive influence on the performance of downstream tasks in both ProtST and PAAG. Additionally, the performance improvement of PAAG surpasses that of ProtST in 10 out of 14 cases, further validating the effectiveness of PAAG in augmenting the quality of protein representation.
4.3 Unconditional Protein Generation
@@
  • Distinct-n Li et al. (2016) is a classical metric in natural language processing that measures textual diversity of generated text by counting distinct n-grams. We use this metric to assesses the fraction of repetitive sequence motifs in sequences from $\mathcal{S}$, which exhibits the biological importance Andrade et al. (2000). A higher Distinct-n suggests fewer repetitive amino acid segments. We set $n=2$ here.
  • Diversity measures the dissimilarity of sequences in $\mathcal{S}$. We employ Mmseq2 Steinegger and Söding (2017) to compute the dissimilarity between each pair of sequences, and utilize the mean of these dissimilarities as Diversity. A higher Diversity signifies a greater diversity in $\mathcal{S}$.
  • Novelty measures the novelty of sequences in $\mathcal{S}$ compared to a reference set. We take UniprotKB  Consortium (2022) as the reference set. For each sequences in $\mathcal{S}$, we employ Mmseq2 Steinegger and Söding (2017) to return the dissimilarity score of the most similar sequence in UniprotKB. Novelty is defined as the mean of dissimilarity score for all sequences in $\mathcal{S}$. A higher Novelty indicates the generated sequences exhibit substantial novelty comparing with the reference set.
Table 2: Results of the unconditional generation task. The values in parentheses $\Delta\%$ represent the absolute relative difference from the values of Natural.
Model Distinct-n ($\Delta\%$) Diversity($\Delta\%$) Novelty $\uparrow$
Natural 235.76(0) 0.829(0) -
@@
ProGEN 165.54(29.78%) 0.845(1.93%) 0.374
Chroma 175.72(25.47%) 0.855(3.14%) 0.638
PAAG 231.88(1.66%) 0.815(1.69%) 0.766
@@
  • PAAG exhibits the closet Distince-n score compared with natural proteins, which proves that proteins generated by PAAG possess similar amino acid distribution as natural proteins.
  • PAAG also maintains the closest diversity to natural proteins, proving PAAG has the similar distribution at protein level.
  • ProGen and Chroma have considerably lower Distince-n score in comparison to PAAG and natural proteins, implying an abundance of repetitions within their generated protein sequence.
In a summary, PAAG is capable of generating high-quality protein sequences, aided by the alignment process.
Refer to caption Refer to caption
Figure 4: Figure (a) and (b) show the $\text{SR}_{e}$ on zinc-finger domain and immunoglobulin domain over all models. Figure (c) and (d) show their distributions of e-value. White bar indicates the mean e-value of each set. PAAG consistently exhibits better performance on all metrics compared with other models. Fine-tuning also introduces additional improvement for PAAG.
Figure 5: Visualization of the generated results on zinc-finger and immunoglobulin domain. The corresponding prompt and generation qualify (e-value) is listed below.
4.4 Protein Design with Domain Annotations
@@
Settings: We utilize two biologically significant domains, zinc-finger domain Klug and Rhodes (1987) and immunoglobulin domain Brummendorf (1995), as the target domain annotation to generate the proteins respectively. For each case, we generate $N=300$ protein sequences given the length of the proteins and the annotation set containing the domain annotations of “zinc-finger” or “immunoglobulin domain”. For all models, the protein sequences are generated with the same length sampled from natural proteins that have corresponding domain in UniprotKB.
@@
Model training: We train PAAG on ProtAnnotation for total 100 epochs, employing a learning rate of 3e-5 and incorporating a warm-up phase with a batch size of 32. Additionally, we extract a subset of ProtAnnotation which only contains proteins with “zinc-finger” and “immunoglobulin domain”. This subset is subsequently utilized to fine-tune the model over 5 epochs, adopting a learning rate of 1e-5 and maintaining the batch size at 32. We denote this fine-tuned version as $\text{PAAG}_{\text{ft}}$. More details are deferred to Appendix A.6.
@@
  • PAAG significantly outperforms the other baselines in success rates $\text{SR}_{1}$ across all tasks by a large margin, e.g. , $51\%$ versus $9\%$ in the immunoglobulin domain task with quality threshold $e=1$. Furthermore, the fine-tuned version $PAAG_{ft}$ can further improve the success rates (54.3% on $\text{SR}_{1}$) consistently, indicating that the finetuning process can help to generate proteins with high quality domains.
  • ProGen outperforms other baseline methods, maintaining consistent results as the threshold changes. This may be due to ProGen remembering the zinc and lg domain sequences present in the training data and output them when using keywords and the first four amino acid sequences as prompts for generation.
@@
Visualizations: We provide the visualization of generated proteins in Figure  5. In Figure 5, all generated domains are marked in red. The textual descriptions generated by the annotation set and the e-value $e$ of this sequence are listed below. The protein structure are folded by Omegafold Wu et al. (2022). As observed in Figure 5, the generated sequences can accurately produce the target domain as per the condition given in the annotation set. Interestingly, in scenarios such as two zinc finger case, when the prompt specifies the presence of multiple zinc finger domains, PAAG generates multiple functional domains in response. However, PAAG falls to capture the precise numbers of these domains, which can be a direction for future improvement.
4.5 Protein Design with Property Annotations
@@
Settings: We employ the subcellular location of proteins as an example property. An additional dataset, termed ProtLocation, is generated by extracting subcellular location labels, encompassing both Bin (2-class) and Sub (10-class) tasks as delineated in Section 4.2, from Deeploc Almagro Armenteros et al. (2017). These labels are then incorporated into annotation-sequence pairs derived from Uniprot. ProtLocation includes $10100$ proteins for training, while a separate set of $2434$ test proteins is reserved for constructing the annotation set for generation. More details can be found in Appendix
Evaluation protocol: For the generated sequences, we employ the official server provided by Deeploc Almagro Armenteros et al. (2017) and Deeploc2 Thumuluri et al. (2022) as the pre-defined oracle to construct the Global-property metric $\mathrm{M}_{T}()$ for predicting binary location label and 10-class label respectively. A generation is deemed successful if the predicted label aligns with the input annotation label.
Results: Table 3 depicts the generation result of PAAG on binary localization annotation. PAAG achieves overall $74.78\%$ success rate, indicting PAAG captures the difference between soluble and membrane-bound, and can generate properties given corresponding property annotations. As for a more challenging setting, PAAG achieves overall $49.71\%$ success rate for the 10-class subcellular localization annotation, still indicating a strong correlation between property annotations and the generated proteins by PAAG. The detailed confusion matrix about 10-class subcellular localization annotation are depicted in Figure 8 in Appendix.
Table 3: The result of protein design with “membrane-bound” and “soluble” annotations.
@@
4.6 Ablation Study

### v3

@@ 3.4. Annotation-guided Protein Design
Definition 3.0 (Annotation-guided Protein Design).
@@
(10) $\vspace{-1.5ex}\max\sum_{T\in\mathcal{A}}\mathrm{M}_{T}(S),\text{s.t. }S=F(\mathcal{A}),$
where $\mathrm{M}_{T}(\cdot)$ is metric function for the annotation $T$.
In this paper, we mainly focus on two type of metrics, functional-domain metric and global-property metric.
  • Functional-domain metric: for the annotation describing a certain functional domain, $\mathrm{M}_{T}()$ will invoke a profile hidden Markov models from Pfam (Mistry et al., 2021) to search for the optimal match within protein $S$ regarding to the domain described by annotation $T$ and assign an e-value $s$ to this match. Given an e-value threshold $e$, we have:
    (11) $\vspace{-2ex}\mathrm{M}_{T,e}(S)=\mathds{1}_{S}(s<e),$
@@ 4. Experiment
4.1. Construction of ProtAnnotation Dataset
To enable multi-level alignment, we build the ProtAnnotation dataset with annotation-sequence pair set for protein design task. Specifically, we select the proteins with “Domain” entry in UniProtKB (Consortium, 2022) to build ProtAnnotation dataset.
However, since the biases within our training dataset will significantly impact the performance of PAAG, we limit our selection to protein-annotation pairs from a refined subset of UniProtKB, namely Swiss-Prot. Each entry in Swiss-Prot has been manually reviewed and supplemented with detailed information with protein function, structure, and interactions. Ultimately, the ProtAnnotation consists of 129,727 proteins. Moreover, in our alignment, we incorporate temperature (Wu et al., 2018) to control the similarity scores of distribution between positive and negative pairs. An appropriate temperature can soften these scores, making the model less sensitive to noisy samples. This adaptability allows the learning process to focus more on meaningful correlations and reduce the impact of irrelevant variations, ultimately enhancing the PAAG’s robustness.
The domain annotations are extracted by these “Domain” entries including the domain description and start & end index of this domain. Additionally, we select four properties as the property annotations, that is, “protein_name”, “organism_name”, “length” and “SIMLARITY”. The textual description is assembled by the template function $G$, which can accept any subset of annotation as input. More details, including the data examples are deferred in Appendix A.2.
4.2. Quality of Aligned Representation
We first conduct multiple experiments on predictive tasks, to evaluate the quality of protein representation produced by PAAG.
Settings: To make a fair comparison with ProtST, we use the same proteins in dataset ProtDescribe (Xu et al., 2023) to first pretrain PAAG, followed by full-model fine-tuning on various downstream tasks. For the full-model fine-tuning, we add a task head for each task and fine-tune the model for 100 epochs. We use the validation set to select the model and report the results on random seed 0, adhering to the same settings as in ProtST (Xu et al., 2023). More details are deferred in Appendix A.6.
@@ 4.2. Quality of Aligned Representation
  • Protein Localization Prediction aims to forecast the subcellular locations of proteins. Derived from DeepLoc (Almagro Armenteros et al., 2017), we focus on two similar tasks. The subcellular localization prediction (Abbr. as, Sub) encompasses 10 location categories and binary localization prediction (Abbr. as, Bin) that includes 2 location categories, soluble and membrane-bound. The splits of data follow the original split in DeepLoc.
  • Fitness Landscape Prediction is primarily focused on the prediction of the effects of residue mutations on the fitness of proteins. We evaluate our models on $\beta$-lactamase (Abbr., $\beta$-lac) landscape from PEER (Xu et al., 2022), the AAV and Thermostability (Abbr., Thermo) landscapes from FLIP (Dallago et al., 2021), and the Fluorescence (Abbr., Flu) and Stability (Abbr., Sta) landscapes from TAPE (Rao et al., 2019). The splits of data follow the splitting setting of ProtST (Xu et al., 2023)
Baselines: We adopt two types of baseline. The first type is the models trained from scratch, including CNN (Shanehsazzadeh et al., 2020), ResNet (Rao et al., 2019), LSTM (Rao et al., 2019), Transformer (Rao et al., 2019). The second type is the pre-trained models with full model tuning, including OntoProtein (Zhang et al., 2022), ProtBert (Elnaggar et al., 2021) and ESM2 (Lin et al., 2023). For the ProtST and PAAG, we train two variants with different initialization weights from ProtBert and ESM2 to verify the enhancement of text knowledge on protein representations from different protein encoders. We utilize distinct subscripts to denote the initialized parameters, such as $\text{PAAG}_{\text{ProtBert}}$ and $\text{PAAG}_{\text{ESM2}}$.
Table 1. Results on protein localization prediction and protein landscape prediction benchmarks. Bold in green and underlined numbers indicate the best and the second best result, respectively.
@@
$\textbf{ProtST}_{\text{ProtBert}}$ 91.78 78.71 0.863 0.804 0.673 0.679 0.745
$\textbf{ProtST}_{\text{ESM2}}$ 92.52 80.22 0.879 0.825 0.682 0.682 0.738
$\textbf{PAAG}_{\text{ProtBert}}$ 92.63 78.96 0.820 0.825 0.668 0.680 0.788
$\textbf{PAAG}_{\text{ESM2}}$ 92.46 81.30 0.888 0.839 0.684 0.682 0.737
Results: Table 1 reports the results of all models on seven baselines. As illustrated in Table 1, PAAG achieves superior performance in comparison to the baselines on all 7 tasks. These results highlight the robust generalization capabilities of PAAG for downstream tasks. Specifically, PAAG outperforms the vanilla pretrained models, ProtBert and ESM2, in all cases, indicating that PAAG can further enhance the quality of protein representation by incorporating the knowledge from textual annotations. According to the results in Table 1, multi-modal training has a positive influence on the performance of downstream tasks in both ProtST and PAAG. Additionally, the performance improvement of PAAG surpasses that of ProtST in 10 out of 14 cases, further validating the effectiveness of PAAG in augmenting the quality of protein representation.
4.3. Unconditional Protein Generation
@@
  • Distinct-n (Li et al., 2016) is a classical metric in natural language processing that measures textual diversity of generated text by counting distinct n-grams. We use normalized Distinct-n to assess the fraction of repetitive sequence motifs in sequences from $\mathcal{S}$, which exhibits the biological importance (Andrade et al., 2000). A higher Distinct-n suggests fewer repetitive amino acid segments. We set $n=2$ here.
  • Diversity measures the dissimilarity of sequences in $\mathcal{S}$. We employ Mmseq2 (Steinegger and Söding, 2017) to compute the dissimilarity between each pair of sequences, and utilize the mean of these dissimilarities as Diversity. A higher Diversity signifies a greater diversity in $\mathcal{S}$.
  • Novelty measures the novelty of sequences in $\mathcal{S}$ compared to a reference set. We take UniprotKB  (Consortium, 2022) as the reference set. For each sequences in $\mathcal{S}$, we employ Mmseq2 (Steinegger and Söding, 2017) to return the dissimilarity score of the most similar sequence in UniprotKB. Novelty is defined as the mean of dissimilarity score for all sequences in $\mathcal{S}$. A higher Novelty indicates the generated sequences exhibit substantial novelty comparing with the reference set.
Table 2. Results of the unconditional generation task. The values in parentheses $\Delta\%$ represent the absolute relative difference from the values of Natural.
Model Distinct-2 ($\Delta\%$) Diversity($\Delta\%$) Novelty $\uparrow$
Natural 0.4309(0) 0.829(0) -
@@
ProGEN 0.3003(30.31%) 0.845(1.93%) 0.374
Chroma 0.3211(25.48%) 0.855(3.14%) 0.638
ProteinDT 0.4909(13.92%) 0.814(1.81%) 0.578
PAAG 0.4314(0.12%) 0.815(1.69%) 0.766
Refer to caption Refer to caption
Figure 3. Figure (a) and (b) show the $\text{SR}_{e}$ on zinc-finger domain and immunoglobulin domain over all models. Figure (c) and (d) show their distributions of e-value. White bar indicates the mean e-value of each set. PAAG consistently exhibits better performance on all metrics compared with other models. Fine-tuning also introduces additional improvement for PAAG.
Figure 4. Visualization of the generated results on zinc-finger and immunoglobulin domain. The corresponding prompt and generation qualify (e-value) is listed below.
@@
  • PAAG exhibits the closest Distince-n score compared with natural proteins, which proves that proteins generated by PAAG possess similar amino acid distribution as natural proteins.
[…]
  • ProGen and Chroma have considerably lower Distince-n score in comparison to PAAG and natural proteins, implying an abundance of repetitions within generated protein sequence. While ProteinDT has higher Distince-n score, it also fails to capture the intrinsic amino acid distrbution in natrual proteins.
In a summary, PAAG is capable of generating high-quality protein sequences, aided by the the multi-level alignment process.
4.4. Protein Design with Domain Annotations
@@
Settings: We utilize two biologically significant domains, zinc-finger domain (Klug and Rhodes, 1987) and immunoglobulin domain (Brummendorf, 1995), as the target domain annotation to generate the proteins respectively. For each case, we generate $N=300$ protein sequences given the length of the proteins and the annotation set containing the domain annotations of “zinc-finger” or “immunoglobulin domain”. For all models, the protein sequences are generated with the same length sampled from natural proteins that have corresponding domain in UniprotKB. To further evaluate the generalization ability of PAAG, we further test our model on generating proteins with EGF-like domain. The details can be found in Appendix A.7.1.
@@
Model training: We train PAAG on ProtAnnotation for total 100 epochs, employing a learning rate of 3e-5 and incorporating a warm-up phase with a batch size of 32. Additionally, we extract a subset of ProtAnnotation which only contains proteins with “zinc-finger” and “immunoglobulin domain”. This subset is subsequently utilized to fine-tune the model over 5 epochs, adopting a learning rate of 1e-5 and maintaining the batch size at 32. We denote this fine-tuned version as $\text{PAAG}_{\text{ft}}$, and the PAAG without fine-tuning is denoted as PAAG. The generalization ability of $\text{PAAG}_{\text{ft}}$ is also evaluated in  A.4. More details are deferred to Appendix A.7.2.
@@
  • PAAG achieves significantly higher success rates $\text{SR}_{1}$ across all tasks by a large margin, e.g. , $51\%$ versus $22\%$ in immunoglobulin domain task with quality threshold $e=1$. Furthermore, the fine-tuned $PAAG_{ft}$ can further improve the success rates (54.3% on $\text{SR}_{1}$) consistently, indicating the finetuning process can help to further improve the quality.
  • ProGen and ProteinDT outperform other baseline methods. This may be due to they memorizing proteins with zinc and lg domain in the training data and output them when using keywords or textual prompts.
@@
Visualizations: We provide the visualization of generated proteins, folded by Omegafold (Wu et al., 2022), in Figure  4. The figure  4 highlights generated domains in red and provides textual descriptions and e-value $e$ for each sequence. We observe the generated sequences accurately produce target domains as specified in the annotation set. Interestingly, in scenarios such as two zinc finger cases, when the prompt specifies the presence of multiple zinc finger domains, PAAG generates multiple functional domains in response. However, PAAG fails to capture the precise numbers of these domains, which can be a direction for future improvement.
Prompt: We further investigate the relationship between the number of domains in prompts and the proteins generated by PAAG. By varying only the domain count in prompts, we generate 900 proteins with ‘Small (1-3)’, ‘Median (4-6)’, and ‘Large (7-9)’ number of domains. Figure 5 shows that increasing domain numbers in prompts leads to a corresponding rise in generated domains across different e-values, indicating that PAAG can discern and generate the specified number of domains from text prompts through multi-level alignment. Given these observations, another interesting question is whether multiple domains in prompt will increase success rates. To this end, we re-evaluate PAAG using prompts specified only one domain (PAAGsingle ${}_{\text{single domain}}$). The results are in Table 8 in Appendix. We can find while multiple domains improve success rates, PAAGsingle ${}_{\text{single domain}}$ also outperforms the baselines, confirming the importance of multi-level alignment.
Figure 5. The relation of number specified in prompt with generated domains by PAAG.
4.5. Protein Design with Property Annotations
@@
Settings: We employ the subcellular location of proteins as an example property. An additional dataset, termed ProtLocation, is generated by extracting subcellular location labels, encompassing Bin (2-class) task as delineated in Section 4.2, from Deeploc (Almagro Armenteros et al., 2017). These labels are then incorporated into annotation-sequence pairs derived from Uniprot. ProtLocation includes $10100$ proteins for training, while a separate set of $2434$ test proteins is reserved for constructing the annotation set for generation. More details can be found in Appendix A.6.3.
Evaluation protocol: For the generated sequences, we employ the official server provided by Deeploc (Almagro Armenteros et al., 2017) as the pre-defined oracle to construct the Global-property metric $\mathrm{M}_{T}()$ for predicting binary location label. A generation is deemed successful if the predicted label aligns with the input annotation label.
Results: As shown in Table 3, PAAG achieves overall $74.78\%$ success rate, indicting PAAG captures the difference between soluble and membrane-bound, and can generate properties given corresponding property annotations. We next will move to more challenging setting, generating proteins with both domain and property annotations.
Table 3. The result of protein design with “membrane-bound” and “soluble” annotations.
@@
Figure 6. Visualization of jointly generation with domain and property annotations. PAAG is capable of integrating domain and property annotations.
4.6. Case study: Joint Generation with Domain and Property Annotations
In this part, we explore the potential of PAAG in generating proteins guided by the flexible combination of different annotations. Specifically, we generate the zinc-finger proteins in membrane-bound and soluble by the model in Section 4.5 following the same protocol in Section 4.4. Despite this challenging setting, the joint success rate $\text{SR}_{1}$ of generating proteins with zinc-finger and corrected property is $\text{SR}_{1}=10.17\%$, However, given the current success rates of other models in generating zinc-finger proteins have approached zero, the task of producing such proteins with specific domain and property annotations remains unattainable by existing models. Zinc-finger domains are naturally occurring soluble proteins involved in DNA editing. Our generation of zinc-finger proteins anchored to the membrane underscores the efficacy of PAAG in producing novel, non-existent proteins. We showcase four examples of the generated results in Figure  6. As shown in Figure 6, PAAG can successfully generate proteins guided by both domain and property annotations, demonstrating its potential in complex protein design tasks.
4.7. Ablation Study